OpenStudy (anonymous):

At a local produce stand sugar-coated strawberries sell for $3 dollare per pund and chocolate covered bananas sell for $4 per ound. How many punds of strawberries must be purchased to have a 5-pound mixture that costs $3.60 per pound?

OpenStudy (anonymous):

3*x +4*y = 3.6*5 x+y =5

OpenStudy (anonymous):

the answers to choose from are 1, 2, 2.5, and 1.5

OpenStudy (anonymous):

and which one did you arrive at?

OpenStudy (anonymous):

I stink at world problems and Im very confused lol

OpenStudy (anonymous):

solve x+y =5 for y plug that into the other equation.

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

is it 1.5?

OpenStudy (anonymous):

Im so confused -_-

OpenStudy (anonymous):

show me what you tried so far... It's hard to help you based off of 'so confused'.

OpenStudy (anonymous):

I mean am I right? or not? Well I just came up with a anwser 1.5 but idk if thats correct